OpenStudy (anonymous):

Whats the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). please help:)

OpenStudy (sasogeek):

if it is parallel to the line with the given equation, it means they have the same slope. make y the subject of the equation and you'll get the slope. then with the slope you find and the given point, you can pick any other point on the new line say (x,y), and then you can find the equation required...

OpenStudy (anonymous):

that makes sense, thank you

OpenStudy (sasogeek):

use this formula just in case it's not so clear y – y1 = m(x – x1) y - 4 = (3/4)(x- (-3)) solve for y Get more explanation here if it's still unclear http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

that helps a bunch thanks!